Two professors at a local college developed a new teaching curriculum designed to increase students grades in math classes.
Answer #1

b)  np(1-p) =    16*0.55*(1-0.55)=       3.96 ≈ 4.0

d) p value =P(X>12) =1-BINOMDIST(12,16,0.55,TRUE) =  0.028